Calculate the Laplace transform of

β€’ (t) integral from (0 to t) (T)e^(T)dT


1
Expert's answer
2022-06-01T02:50:42-0400
"{\cal L}\{f\}(s)=\int_0^{\infty}f(t)e^{-st}dt""{\cal L}\{t\int_0^tTe^TdT\}(s)=\int_0^{\infty}t\int_0^tTe^TdTe^{-st}dt""=\int_0^{\infty}t(1+e^t(t-1))e^{-st}dt=\frac{3s-1}{s^2(s-1)^3}"